Notice: wiki.ucar.edu has been updated. Please see the announcement for details.

Versions Compared

Key

  • This line was added.
  • This line was removed.
  • Formatting was changed.

...

Removed Licor (sn 1176) from 15m.*  When removing this sensor due to bad chopper motor, it was not running.  Rebooted sensor and heard the chopper motor in the head and it sounded bad.  Removed the 7m Licor (sn????sn0813) and taped up connections and moved this Licor to the 15m height.  Plugged in all connections as normal (signal green cable and power through BNC).  All looked (and sounded) great.  Replaced the TRH fans at heights 7m and 15m.  Both fans sounded REALLY bad.  The 2m TRH fan is did not sounded sound very good either.  Guessing we will be replacing that fan soon. 

...